A cane filled with water is revolved in a vertical circle of radius 4m and the water just does not fall down the time period of revolution will be

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

at the top most position

three forces will act on the water

1 weight in the downward direction

2  normal reaction by bucket in the downward direction

3 centrifugal force in the upward direction

 

 so  mg + N =mv2/r

 for water just does not fall   N ≥ 0

so  mg = mv2/r

 v =√gr

 

 

so time period T = 2∏r/v

                         =2*3.14*4 /√4g

                          =3.14*4 /√9.8

                         =4.01

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

NISA
9 years ago
i did nt understand the solution * this star idicates n my did v take n=o 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

at the top most position
three forces will act on the water
1 weight in the downward direction
2 normal reaction by bucket in the downward direction
3 centrifugal force in the upward direction
so mg + N =mv2/r
for water just does not fall N ≥ 0
so mg = mv2/r
v =√gr
so time period T = 2∏r/v
=2*3.14*4 /√4g
=3.14*4 /√9.8
=4.01
Thanks and Regards